Question 2 equally charged particles, held 4.4 x 10-3 m apart, are released from rest. The initial acceleration of the first particle is observed to be 9.0 m/s? and that of the second 7.8 m/s?. If the mass of the first particle is 8.2 x 107 kg, what are (a) the mass of the second particle and (b) the magnitude of the charge of each particlei (a) Number Units (b) Number Units
Answer #1

Here ,

a) as the force on each of the masses is same

m1 * a1 = m2 * a2

8.2 *10^-7 * 9 = m * 7.8

m = 9.46 *10^-7 Kg

the mass of the second particle is 9.46 *10-7 Kg

b) let the magnitude of charge of each particle is q

k * q^2/d^2 = m1 * a1

9 *10^9 * q^2/(4.4 *10^-3)^2 = 8.2 *10^-7 * 9

solving for q

q = 1.26 *10^-10 C

the charge q is 1.26 *10-10 C